# GraphTeam: Facilitating Large Language Model-based Graph Analysis via Multi-Agent Collaboration

### v2

@@ 3.2. Question Agent
In an undirected graph … The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5).
Q: Is there a path in this graph that visits every node exactly once? …
A: In the answer, …, Output format such as: … 1, 4, 8.
@@
refined_question: Is there a path in this graph that visits …
@@
input_data: ’Nodes’: [0, 1, 2, 3, 4, 5, 6],
’Edges’: [[0, 3], [0, 1], [1, 4], [1, 6], [2, 4], [3, 5], [3, 6], [4, 5]]
@@
3.3. Search Agent
During the solving process of a hard problem, it is common for human beings to seek help from external sources (Eisenberg and Berkowitz, 1990) or make analogy to a previously solved one (Bybee et al., 2006). Therefore, we first build a knowledge base with documentation and experience information to enhance the programming ability of the multi-agent system, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
3.3.1. Knowledge Base Construction
@@
description: Returns a Hamiltonian path in the given …
parameters: ”G : NetworkX graph” , …
returns: ”path : list”: ”A list of nodes which form a Hamiltonian path in G.”
examples:
    import NetworkX as nx
    nx.tournament.hamiltonian_path(G)
@@
Experience. We use the system without experience base as the solver, and apply it on the training set to collect experiences. Note that each problem in the training set $\mathcal{D}_{\text{train}}$ is usually associated with problem type $t\in\mathcal{T}$. Thus we collect candidate experiences $\mathcal{K}_{exp}^{t}$ for each problem type $t$, and evaluate the utility of each experience on the validation set $\mathcal{D}_{\text{valid}}$. Only the experience with the highest utility will be kept in each $\mathcal{K}_{exp}^{t}$.
[…]
Experience Information
question: In an undirected graph, …, The nodes are numbered from 0 to 6, and the edges are: (0,3) … (4,5) Q: Is there a path in this graph …
answer: Yes. The path can be: 0, 1, 6, 3, 5, 4, 2.
meta information:
thought: The experience from this task involved solving the problem of finding a Hamiltonian path. …
code:
    import NetworkX as nx
    def hamiltonian_path_backtrack(graph, path):
3.3.2. Knowledge Retrieval
@@
For effective information retrieval, we build vector databases with LlamaIndex (lla, 2024) for documentation and experiences, respectively. Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database. Note that the experiences of similar problems contains detailed solutions, and are usually more helpful for downstream agents than documentations. Therefore, we first use similarity matching via LlamaIndex to score the experiences. If the largest similarity score exceeds a predefined threshold $\delta$, the corresponding experience will be directly returned by the search agent without querying related documentation information. Otherwise, the search agent will return the most related entries in documentation base. Finally, either the experience or the documentation information returned by the search agent will be passed to the downstream agents for retrieval-augmented generation (RAG). For simplicity, we denote the operation of the search agent as
(2) $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta).$
Algorithm 1 Problem-Solving Experience Collection
1:Training dataset $\mathcal{D}_{\text{train}}$, validation dataset $\mathcal{D}_{\text{valid}}$, candidate experience size $N_{exp}$ for each problem type $t\in\mathcal{T}$.
2:Experience knowledge base $\mathcal{K}_{exp}$.
3:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{train}}$ do
4:     if $|\mathcal{K}_{exp}^{t}|$ $N_{exp}$ then
[…]
6:         if answer equals to ground truth $a$ then
[…]
8:         end if
9:     end if
10:end for
11:for each problem $(q,a,t)$ from $\mathcal{D}_{\text{valid}}$ do
12:     for $e\in\mathcal{K}_{exp}^{t}$ do
[…]
14:         if answer equals to ground truth $a$ then
15:              $\text{Utility}_{e}\leftarrow\text{Utility}_{e}+1$
16:         end if
17:     end for
18:end for
19:for $t\in\mathcal{T}$ do
20:     $\mathcal{K}_{exp}^{t}\leftarrow\text{argmax}_{e\in\mathcal{K}_{exp}^{t}}\text{Utility}_{e}$
21:end for
22:$\mathcal{K}_{exp}\leftarrow\{\mathcal{K}_{exp}^{t}\}_{t\in\mathcal{T}}$
3.4. Coding and Reasoning Agents
@@
Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems. In practice, the generated codes may encounter compilation or running errors. Following the trial-and-error strategy (Robson and Pólya, 1945; Ramnarine-Rieks, 2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages. More formally, the operation of coding agent in the $n$-th trial can be written as:
@@
(3) $\begin{split}(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},\text{knowledge}_{q},\\ \{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1}),\end{split}$
@@
If the coding agent fails to fix the codes before reaching the maximum number of trials, we will employ the reasoning agent to directly answer the question without programming:
(4) $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f}).$
3.5. Answer Agent
Based on the output format $q_{f}$, the answer agent will organize the results given by the coding or reasoning agents as final results. Similar to the retry mechanism of coding agent, we also introduce a self-checking mechanism that iteratively refines the results according to format requirement $q_{f}$. Formally, the operation of answer agent can be written as:
(5) $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1}),$
@@ 3.5. Answer Agent
3.6. Summary
We formalize the inference pipeline of GraphTeam in Alg. 2. Most agent functions (including Question, Coding, Reasoning and Answer) in Alg. 2 are all implemented by an LLM with different prompt templates. The Search function is implemented based on LlamaIndex (lla, 2024), and the Solver function in Alg. 1 is implemented by a simplified version of GraphTeam without experience base. We will provide more implementation details in the appendix.
@@ 3.6. Summary
1:Problem described by question $q$, knowledge base $\mathcal{K}$, threshold $\delta$ for determining whether experience or documentation is retrieved, maximum number of retries $N_{retry}$, number of self-checking iterations $N_{check}$.
2:Final answer of the problem $\text{output}_{q}$.
3:$(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
4:$\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
5:$n=0$
6:while $\text{error}_{q}^{n}\neq\tt{None}$ and $n<N_{retry}$ do
7:     ++n
@@
8:     $(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},$ $\text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1})$
9:     $\text{result}_{q}\leftarrow\text{result}_{q}^{n}$
10:end while
11:if $\text{error}_{q}^{n}\neq\tt{None}$ then
12:     $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f})$
13:end if
14:$\text{output}_{q}^{0}\leftarrow\text{result}_{q}$
15:for $n=1,2,\dots,N_{check}$ do
16:     $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1})$
17:end for
4. Experiments
In this section, we conduct experimental evaluation to answer the following research questions (RQs): RQ1: How effective is our proposed GraphTeam compared to the state-of-the-art (SOTA) baselines? RQ2: Have each component of GraphTeam played their roles effectively? RQ3: How does our framework perform with respect to different task categories and output formats of graph analysis problems? RQ4: How do the hyper-parameters of certain components in the system have an impact on the overall performance?
4.1. Experimental Setup
4.1.1. Datasets
@@
To evaluate the effectiveness of our system in solving graph analysis problems, we use six benchmarks for LLM-based graph analysis: Talk like a Graph (Fatemi et al., 2023) primarily studies some fundamental problems in graph analysis and explores the cognitive abilities of LLMs regarding graphs. LLM4DyG (Zhang et al., 2024b) focuses on basic problems related to dynamic graphs, while GraphWiz (Chen et al., 2024a), NLGraph (Wang et al., 2024a), and Graph Instruct (Luo et al., 2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory. GNN-AutoGL (details shown in Appendix B) is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL (Guan et al., 2021).
4.1.2. Metrics
@@
For benchmarks with definite answers (Fatemi et al., 2023; Chen et al., 2024a; Wang et al., 2024a; Luo et al., 2024; Zhang et al., 2024b), we use accuracy as the evaluation metric. For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria.
Specifically, we extract the generated code from the model response and compare the execution result with the reference solution to assess whether the model-generated results match the given standard answers or reference solutions. For the first five benchmarks (Fatemi et al., 2023; Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024; Wang et al., 2024a), the evaluation results are binary based on the exact match between the execution results and the standard answers or reference solutions. For GNN-AutoGL, we first evaluate whether the code can be executed without errors; if it cannot, it scores 0. Otherwise, we check whether the hyper-parameters(eg. model name, dataset name, hidden dimension, dropout) meet the requirements. Each correct hyper-parameter earns partial credit, and if all hyper-parameters are correct, the score is 1.
4.1.3. Settings
@@
We choose GPT-4o-mini (OpenAI, 2024) as the base LLM for all agents of our GraphTeam. For Python libraries, we use NetworkX version 3.3111https://networkx.org/ and AutoGL version 0.4.0222https://mn.cs.tsinghua.edu.cn/autogl/documentation/, together with Python version 3.10.14. The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$. The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
4.1.4. Baselines
@@
As the SOTA methods for each benchmark are quite diverse and some of which are specialized for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM (Ouyang et al., 2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers (Sanford et al., 2024) as the SOTA; for GraphWiz, LLM4DyG, and GraphInstruct (Chen et al., 2024a; Zhang et al., 2024b; Luo et al., 2024), we select the best model in their benchmark papers. For GraphWiz (Chen et al., 2024a), we did not choose GraphAgent-Reasoner (Hu et al., 2024) because it only provide results for 6 out of 9 categories.
4.2. Main Results (RQ1)
@@
Table 1. Performance comparison on six graph analysis benchmarks in terms of accuracy (%).
@@
Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
Main GPT-4o-mini 71.67 47.61 51.40 58.03 53.67 3.28
SOTA 97.30 65.00 79.70 53.71 43.53 -
GraphTeam 99.13 88.61 97.80 94.81 88.14 48.44
     - question agent 85.87 62.81 79.22 59.18 85.86 14.53
Input-Output Normalization      - answer agent 94.80 87.97 96.70 94.27 85.25 45.78
     - documentation 92.03 87.97 96.40 94.07 86.12 46.88
Knowledge Retrieval      - experience 89.96 82.83 54.70 60.26 76.64 4.76
     - coding agent 69.40 46.39 39.30 59.33 61.09 -
     - retry 91.90 83.53 87.70 89.33 85.66 44.22
Problem Solving      - reasoning agent 92.77 54.17 96.90 94.00 85.45 46.36
Table 2. Performance comparison on the hard subset of six graph analysis benchmarks in terms of accuracy (%).
Category Models Talk like a Graph GraphWiz NLGraph LLM4DyG GraphInstruct GNN-AutoGL
Main GPT-4o-mini 47.90 28.92 44.93 42.56 23.19 3.28
SOTA 93.90 70.50 44.96 33.90 11.71 -
GraphTeam 99.20 76.25 98.14 98.33 73.33 48.44
     - question agent 70.00 39.75 69.82 53.67 73.29 14.53
Input-Output Normalization      - answer agent 97.40 75.25 97.51 96.33 70.00 45.78
     - documentation 98.90 75.25 97.51 94.33 69.76 46.88
Knowledge Retrieval      - experience 98.40 62.92 42.32 70.56 63.95 4.76
     - coding agent 64.00 38.33 19.21 51.56 38.24 -
     - retry 95.40 61.00 40.27 85.44 70.95 44.22
Problem Solving      - reasoning agent 98.90 66.08 97.51 95.67 69.62 46.36
We present the results in Tab.1. GraphTeam surpasses previous SOTA by 1.83-44.61% in accuracy, and outperforms the base LLM (GPT-4o-mini) by 27.46-46.40% in accuracy. Such significant improvements fully demonstrate the effectiveness of our carefully designed multi-agent system. Besides, the time cost for answering each problem is less than 3 seconds, and the monetary cost of API usage is less than \$0.01.
To further demonstrate GraphTeam’s performance on challenging problems, we assess its performance on the “hard” subset of each benchmark. For GraphWiz, we adhere to the official partition and report results on their designated hard set. For other benchmarks, which all have their problems categorized into different types or domains, we employ a model-based approach: we identify the three categories where previous SOTA performs poorest and select problems from these categories. This selection method ensures we target the most challenging aspects of each benchmark.
Tab. 2 summarizes the system’s performance on these high-difficulty questions. Notably, GraphTeam maintains high accuracy on the Talk like a Graph hard set (99.20%), and shows significant improvements over SOTA for NLGraph (98.14% v.s. 44.96%) and LLM4DyG (98.33% v.s. 33.90%). While performance drops are observed for GraphWiz and GraphInstruct compared to the general set, GraphTeam still substantially outperforms both GPT-4o-mini and SOTA methods. This consistent and strong performance across diverse hard sets underscores GraphTeam’s robustness and its ability to effectively handle complex graph-related tasks.
4.3. Ablation Study (RQ2)

### v3

@@ 3.2 Question Agent
In an undirected graph …, the edges are: (0,3) … (4,5).
Q: Is there a path … visits every node exactly once? …
A: In the answer, …, Output format such as : …
@@
refined_question: Is there a path in this graph that …
@@
input_data: Nodes: [0, 1, 2, 3, 4, 5, 6],
Edges: [0, 3], [0, 1], [1, 6], [2, 4], [3, 5], [3, 6], [4, 5]
@@
3.3 Search Agent
During the solving process of a hard problem, it is common for human beings to seek help from external sources Eisenberg and Berkowitz (1990) or make analogy to a previously solved one Bybee et al. (2006). Therefore, we first build a knowledge base with documentation and experience information, and then ask the search agent to retrieve relevant entries from the knowledge base for each question.
3.3.1 Knowledge Base Construction
@@
description: Returns a Hamiltonian path …
parameters: G : NetworkX graph, …
returns: path: list…
examples: …
Experience Information
question: In an undirected graph, …, the edges are: (0,1) … (5,6) Q: Is there a path in this graph …
answer: Yes. The path can be: 0, 1, 6, 5.
thought: This is finding a Hamiltonian path. …
code:
    def hamiltonian_path(graph, path):
@@
Experience. We use the system without experience base as the solver, and apply it on the training set to collect experiences. Note that each problem in the training set $\mathcal{D}_{\text{train}}$ is usually associated with problem type $t\in\mathcal{T}$. Thus we collect candidate experiences $\mathcal{K}_{exp}^{t}$ for each problem type $t$, and evaluate the utility of each experience on the validation set $\mathcal{D}_{\text{valid}}$. Only the experience with the highest utility will be kept in each $\mathcal{K}_{exp}^{t}$. More details including the pseudo codes for building the experience base can be found at Alg. 2 in Appendix A.1.
3.3.2 Knowledge Retrieval
@@
For effective information retrieval, we build vector databases with LlamaIndex Liu (2022) for documentation and experiences, respectively. Then the search agent will take the refined question $q_{r}$ and graph type $q_{t}$ as query, and find the most relevant experience and API information from the database. Note that the experiences of similar problems contains detailed solutions, and are usually more helpful for downstream agents than documentations. Therefore, we first use similarity matching via LlamaIndex to score the experiences. If the largest similarity score exceeds a predefined threshold $\delta$, the corresponding experience will be directly returned by the search agent without querying related documentation information. Otherwise, the search agent will return the most related entries in documentation base. Finally, either the experience or the documentation information returned by the search agent will be passed to the downstream agents for retrieval-augmented generation (RAG). For simplicity, we denote the operation of the search agent as
$\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta).$ (2)
3.4 Coding and Reasoning Agents
@@
Given the outputs of question and search agents, the coding agent will generate and execute Python codes to solve the problems. In practice, the generated codes may encounter compilation or running errors. Following the trial-and-error strategy Robson and Pólya (1945); Ramnarine-Rieks (2011) for problem solving, we introduce a retry mechanism that attempts to fix the codes with previous codes and error messages. More formally, the operation of coding agent in the $n$-th trial can be written as:
@@
$\begin{split}(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},\\ \text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1}),\end{split}$ (3)
@@
If the coding agent fails to fix the codes when reaching the maximum number of trials, we will employ the reasoning agent to directly answer the question without programming:
$\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f}).$ (4)
In this way, the coding and reasoning agents can solve problems collectively, and achieve a collaborative functionality at the system level.
3.5 Answer Agent
Based on the output format $q_{f}$, the answer agent will organize the results given by the coding or reasoning agents as final results. Similar to the retry mechanism of coding agent, we also introduce a self-checking mechanism that iteratively refines the results according to format requirement $q_{f}$. Formally, the operation of answer agent is:
$\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1}),$ (5)
@@ 3.5 Answer Agent
3.6 Summary
We formalize the inference pipeline of GraphTeam in Alg. 1. Most agent functions (including Question, Coding, Reasoning and Answer) in Alg. 1 are all implemented by an LLM with different prompt templates. The Search function is implemented based on LlamaIndex Liu (2022). We will provide more implementation details in Appendix A.
@@ 3.6 Summary
0:  Problem described by question $q$, knowledge base $\mathcal{K}$, threshold $\delta$ for determining whether experience or documentation is retrieved, maximum number of retries $N_{retry}$, number of self-checking iterations $N_{check}$.
0:  Final answer of the problem $\text{output}_{q}$.
1:  $(q_{r},q_{t},q_{g},q_{f})\leftarrow\operatorname{Question}(q)$
2:  $\text{knowledge}_{q}\leftarrow\operatorname{Search}(q_{r},q_{t},\mathcal{K},\delta)$
3:  $n=0$
4:  while $\text{error}_{q}^{n}\neq\tt{None}$ and $n<N_{retry}$ do
5:     ++n
@@
6:     $(\text{result}_{q}^{n},\text{code}_{q}^{n},\text{error}_{q}^{n})\leftarrow\operatorname{Coding}(q_{r},q_{t},q_{g},q_{f},$ $\text{knowledge}_{q},\{\text{code}_{q}^{i}\}_{i=1}^{n-1},\{\text{error}_{q}^{i}\}_{i=1}^{n-1})$
7:     $\text{result}_{q}\leftarrow\text{result}_{q}^{n}$
8:  end while
9:  if $\text{error}_{q}^{n}\neq\tt{None}$ then
10:     $\text{result}_{q}\leftarrow\operatorname{Reasoning}(q_{r},q_{t},q_{g},q_{f})$
11:  end if
12:  $\text{output}_{q}^{0}\leftarrow\text{result}_{q}$
13:  for $n=1,2,\dots,N_{check}$ do
14:     $\text{output}_{q}^{n}\leftarrow\operatorname{Answer}(q_{f},\text{output}_{q}^{n-1})$
15:  end for
4 Experiments
We conduct extensive experiments on six graph reasoning benchmarks to validate the effectiveness of GraphTeam.
4.1 Experimental Setup
4.1.1 Datasets
@@
To evaluate the effectiveness of our system, we consider six benchmarks for LLM-based graph reasoning: Talk like a Graph Fatemi et al. (2024) primarily studies some fundamental problems in graph reasoning and explores the cognitive abilities of LLMs regarding graphs. LLM4DyG Zhang et al. (2024b) focuses on basic problems related to dynamic graphs, while GraphWiz Chen et al. (2024a), NLGraph Wang et al. (2024a), and GraphInstruct Luo et al. (2024) are used to evaluate the understanding and mastery of algorithms in fundamental graph theory. GNN-AutoGL is a benchmark we constructed in this paper to study the ability of LLMs in deploying GNNs with AutoGL Guan et al. (2021). Dataset statistics and details can be found at Appendix B.1 and  B.2.
4.1.2 Metrics
@@
For benchmarks with definite answers Fatemi et al. (2024); Wang et al. (2024a); Zhang et al. (2024b); Chen et al. (2024a); Luo et al. (2024), we use accuracy as the evaluation metric. For GNN-AutoGL, since the task requires a complex pipeline (including identifying proper APIs, filling correct hyper-parameters, successfully training and evaluating the corresponding GNN model on a specified dataset, we evaluate whether the generated codes can be executed correctly and whether the key hyper-parameters (e.g., dataset name, model name, hidden dimension, dropout) are correct as the evaluation criteria. More details are shown in Appendix B.3.
4.1.3 Settings
@@
We choose GPT-4o-mini Hurst et al. (2024) and Qwen2.5-Coder-7B Hui et al. (2024) as the base LLM for the agents of GraphTeam. For Python libraries, we use NetworkX version 3.3 Hagberg et al. (2008) and AutoGL version 0.4.0 Guan et al. (2021), together with Python version 3.10.14. The hyper-parameters are set as follows: the number of candidate experiences $N_{exp}=10$, similarity matching threshold for using experience $\delta=0.85$. The maximum number of retries for the coding agent is $3$, and the number of self-checking iterations for the answer agent is $3$.
4.1.4 Baselines
@@
As the SOTA methods for each benchmark are quite diverse and some of which are designed for a specific benchmark, we directly compare with the previously reported SOTA performance for each benchmark: for NLGraph, we select GUNDAM Ouyang et al. (2024) as the SOTA; for Talk like a Graph, we select the fine-tuned transformers Sanford et al. (2024) as the SOTA; for LLM4DyG, GraphWiz, and GraphInstruct Zhang et al. (2024b); Chen et al. (2024a); Luo et al. (2024), we select the best model in their benchmark papers. In addition, we adopt AutoGPT Significant Gravitas (2023) and Open Interpreter Open Interpreter (2023) as single-agent baselines, while employing AutoGen Wu et al. (2024) and ChatDev Qian et al. (2024) as multi-agent baselines for comparison.
4.2 Main Results
@@
Table 1: Performance comparison between GraphTeam and baselines in terms of accuracy (%).
@@
Base LLM       Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL Average Rank
-       Previous SOTA 79.7 90.2 43.5 65.0 53.7 - 5.5
      Base Model 51.4 61.3 53.7 47.6 58.0 3.3 7.5
      AutoGPT 65.8 65.2 58.5 50.2 61.6 10.3 5.7
      Open Interpreter 68.2 85.6 54.2 53.9 55.5 8.8 5.5
      AutoGen 59.7 70.2 57.7 62.8 47.3 28.1 5.8
      ChatDev 79.6 72.5 70.6 67.8 77.8 34.5 3.5
GPT-4o-mini       GraphTeam 97.8 94.5 88.1 88.6 94.8 48.4 1.0
      Base Model 37.3 63.1 34.5 51.8 38.2 1.9 8.3
Qwen2.5-Coder-7B       GraphTeam 93.0 85.6 84.5 78.6 87.4 36.7 2.2
As shown in Table 1, GraphTeam achieves significant and consistent improvements over baseline methods on all six benchmarks. Compared with the best baseline, GraphTeam (GPT-4o-mini) has 15.3% accuracy gains on average. Even when equipped with open-source model, GraphTeam (Qwen2.5-Coder-7B) still has 7.5% average accuracy improvements over the best baseline based on GPTs. These results demonstrate that GraphTeam can achieve promising performance with both open-source and closed-source LLMs. Among the baselines, LLM-based agents (i.e., AutoGPT, Open Interpreter, AutoGen and ChatDev) consistently achieve better performance than the corresponding base LLM (i.e., GPT-4o-mini). For example, the single-agent baseline AutoGPT has 6.1% average accuracy improvements over the base LLM. Moreover, the multi-agent baseline ChatDev achieves even better results, with 21.3% average accuracy gains over the base LLM. These observations support our motivation of employing LLM-based agents for graph reasoning.
4.3 Ablation Study